[Hallc_running] [New Logentry] RC Update - 03/07/22

sjdkay at jlab.org sjdkay at jlab.org
Sun Jul 3 11:15:02 EDT 2022

Logentry Text:
A quick update on the plans today -

Accelerator Update

- Beam was pretty stable through swing and owl, but was limited to 43uA.
- We're currently down and have been since around 07:30.
- The plan from MCC is to just keep delivering as high as they can (probably 43uA) until Tuesday when they will re-assess slightly.

Physics Update

- We finished off the Q2 = 6.00, W = 2.40, x = 0.55, High Epsilon, SHMS Centre LH2 setting - https://logbooks.jlab.org/entry/4010337
- The Aluminium dummy for this setting was also done, we're on the LD2+ version of this setting now. The current projection is that we will hopefully finish this during swing tonight (so long as we get beam back soon).
- Compared to my plan last night, we were only actually ~1hr 30 min behind schedule... but now the beam is down so we're falling a bit further behind that.
- Following discussions last night (https://logbooks.jlab.org/entry/4010192) we will skip some settings and move on to the Fpi points that are more sensitive to the higher beam energy faster.
- This will likely be tweaked again slightly later this evening, I'll update once we know more.

Hall Update

- No issues in the Hall luckily, no more Q1 interlocks since this one yesterday - https://logbooks.jlab.org/entry/4009940


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4010433
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220703/8cee16f5/attachment.html>

More information about the Hallc_running mailing list